OpenStudy (anonymous):

Given: sin(x) = 2/3 and cos(y) = 3/4 Find the values of: a) sin(x+y) b) cos(x-y) c) tan(x+y)

OpenStudy (anonymous):

I can't seem to find x and y

OpenStudy (anonymous):

^Exact values btw

OpenStudy (tommo_lcfc):

Consider 2 different right angled triangles with two of the sides as mentioned in the question.

OpenStudy (anonymous):

How does this give x and y? I need exact values

OpenStudy (tommo_lcfc):

Hang on, I'm working on it.

OpenStudy (anonymous):

inverse functions for cosine and sine

OpenStudy (anonymous):

Yeah, but as I mentioned, I need exact values

OpenStudy (anonymous):

sin(x+y)=sinx*cosy+cox*siny

OpenStudy (anonymous):

how do i get cos x and sin y?

OpenStudy (anonymous):

sinx=2/3 so cosx= sqrt(1-(sinx)^2)

OpenStudy (anonymous):

similarly, siny= sqrt(1- (cosy)^2)

OpenStudy (tommo_lcfc):

sin(x+y) is quite easy. Use the identity that @vedic mentioned. You already know sin(x) and cos(y) from the question. Then you need right angled triangles or follow @vedic's instructions. The first two answers should be 1, the last one should be 8.4. (Please confirm the last one, as I'm unsure.)

OpenStudy (anonymous):

Thanks guys

OpenStudy (anonymous):

I'll work on it

OpenStudy (anonymous):

for tan(x+y).. u cn simply divide sin(x+y) and cos(x+y)

OpenStudy (anonymous):

formulas u cn find frm dis list...

OpenStudy (anonymous):

@tommo_lcfc how u getting answer 1?

OpenStudy (anonymous):

\[\frac{6+\sqrt{35}}{12}\]

OpenStudy (anonymous):

yeaahh thts the answer...

OpenStudy (anonymous):

working on no. 2 now

OpenStudy (anonymous):

\[\frac{3\sqrt{5}+2\sqrt{7}}{12}\]

OpenStudy (anonymous):

no.3 \[\frac{6+\sqrt{35}}{3\sqrt{5}-2\sqrt{7}}\]

OpenStudy (anonymous):

ryt... someone here wants to understand and do things... i appreciate that..:)

OpenStudy (anonymous):

Thanks heaps :D